Reggie is a self-employed taxpayer who turns 59 years old at the end of the year (2018). In 2018, his net Schedule C income was $336,000. This was his only source of income. This year, Reggie is considering setting up a retirement plan. What is the maximum amount he may contribute to the self-employed plan in each of the following situations? a. He sets up a SEP IRA. b. He sets up an individual 401(k).
Answer #1
Particulars SEP IRA 401(k)
a Dollar limit           55,000              55,000
Percent of net income:
b Percentage 20% 20%
c Schedule C net income        3,36,000           3,36,000
Less:
d Deductible SE taxes     11,176.09         11,176.09
e= c-d Net income 3,24,823.91     3,24,823.91
f= b*e Limit on net income     64,964.78         64,964.78
Dollar allowance              17,000
g= minimum of a,f Maximum contribution     55,000.00         55,000.00
h Catch up                 6,000
i= g+h Total contribution     55,000.00         61,000.00